OpenStudy (anonymous):

what is the remainder of (x^2 + 2x + 14) ÷ (x + 7)

OpenStudy (umarsback):

ayyy doing your work hahaaha

OpenStudy (anonymous):

lol yeah ameer's helpin me i hate math,

OpenStudy (anonymous):

use the division algorithm \[x^2 + 2x + 14 = (x+7)q + r\] let x = 7 and you will find your remainder

OpenStudy (anonymous):

-7 sorry

OpenStudy (anonymous):

ok thannk you!(: @jayz657

OpenStudy (anonymous):

your welcome